if z^4= (z-1)^4 then the roots of the equation are represented in the argand plane are by points that are:
a)collinear
b)concyclic
c)vertices of parallelogram
d)none

Grade:Upto college level

1 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Mukul,

z^4=(z-1)^4

z^2=±(z-1)^2

a)let z^2=+(z-1)^2, 2z=1,z=1/2

b)let z^2=-(z-1)^4

2z2-2z+1=0

z=(1/2)±(i/2)

So the solution set consists of 1/2,1/2 +i/2, 1/2 -i/2, which are obviously collinear

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Mukul !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR